OpenStudy (anonymous):

cosx=4/5, csc x<0 please help

OpenStudy (mertsj):

Is there a question here somewhere?

OpenStudy (anonymous):

yes sorry find sin2x, cos2x, tan2x

OpenStudy (anonymous):

It definatley would help to draw a picture! Try that.

OpenStudy (anonymous):

yes you definitely can

OpenStudy (anonymous):

sure you will get a sticker lol

OpenStudy (anonymous):

on no they are different

OpenStudy (anonymous):

ok hold on. let me figure it out

OpenStudy (anonymous):



OpenStudy (anonymous):

right, then mark the sides, I mean opp, hypotenus and adja

OpenStudy (anonymous):

hyp=5 adj=4 we don't have opp

OpenStudy (anonymous):

you are right, so how to find opp?

OpenStudy (anonymous):

hyp^2= adj^2 +opp^2

OpenStudy (anonymous):

yes, yes, yes, pythagorean,!!! give me the answer

OpenStudy (anonymous):

5^2= 4^2+ opp^2

OpenStudy (anonymous):

hold on, what do you want to do? find x ? or find other trigs?

OpenStudy (anonymous):

+or - 3

OpenStudy (anonymous):

find sin 2x, cos 2x and tan 2x

OpenStudy (anonymous):

back to your problem, you have 2 solution is +- 3, choose 1. don't forget, the sign just show where the point is.

OpenStudy (anonymous):

i know that's why I am confused on what to work on since i understand. I posted this four days ago and we worked on it in school.

OpenStudy (anonymous):

choose solution, just 1

OpenStudy (anonymous):

i don't understand what you mean

OpenStudy (anonymous):

you have +- 3, choose 1, 3 or -3?

OpenStudy (anonymous):

I beleive is -3 because csc <0

OpenStudy (anonymous):

right

OpenStudy (anonymous):

yeap

OpenStudy (anonymous):

now, you have opp, can figure out sin and then tan. that's it

OpenStudy (anonymous):

but there is one more problem , you just have cos x , and sinx not cos 2x, or sin 2x. you must convert to those to find the final answer

OpenStudy (anonymous):

sin2x= 2sin xcos x

OpenStudy (anonymous):

yeap.

OpenStudy (anonymous):

-24/5

OpenStudy (anonymous):

cos2x= sin^x- cos^x

OpenStudy (anonymous):

I don't calculate it for you, you must know how to do it. i will not be with you on final test day. ok?

OpenStudy (anonymous):

by the way can i switch the position of sine and cos?

OpenStudy (anonymous):

you have book, and the answer on it. do, and check, if you get stuck, Madaamm, Me here, available for you.. hehehe..

OpenStudy (anonymous):

i know

OpenStudy (anonymous):

on what?

OpenStudy (anonymous):

it is not in the book. it is one of the even #

OpenStudy (anonymous):

it sounds like a pervert is wooing a beaauuutiiifulll woman, right? hahahaha..

OpenStudy (anonymous):

on any of the formulas

OpenStudy (anonymous):

you really sound very happy today when i am all worried about the finals

OpenStudy (anonymous):

surely you can switch the position of a product as ab = ba , don't just replace sin by cos .because you cannot do it

OpenStudy (anonymous):

you forgot the fact that I am qualified enough to not take final, right? my take home final exam is in perfect condition, I just study Cal3 , done, why am I not happy???

OpenStudy (anonymous):

i mean the one for cos2x= sin^2- cos^2

OpenStudy (anonymous):

i mean is the other way around cos^2x - sin^2x

OpenStudy (anonymous):

that's what I am saying.

OpenStudy (anonymous):

NON, NON, Mademoisell, not at all,

OpenStudy (anonymous):

i can't switch?

OpenStudy (anonymous):

whao you are working on your french too? you forgot e at the end.

OpenStudy (anonymous):

no, just a very simple word

OpenStudy (anonymous):

you cannot!! that 's it

OpenStudy (anonymous):

ok tomorrow we will work more on log

OpenStudy (anonymous):

and on simplifying that is really really driving me crazy

OpenStudy (anonymous):

if you stuck at somewhere too anoying, let it there. doesn't matter, go for other question to get credit.

OpenStudy (anonymous):

here is another problem. simplify cos^3 x+ sin^2 xcosx

OpenStudy (anonymous):

factor cos x out

OpenStudy (anonymous):

i did but it does not look right

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

cosx (cos^2 v + sin^2 x) = cos x only, the term in bracket =1

OpenStudy (anonymous):

don't i put +1 next to sin^2x?

OpenStudy (anonymous):

you take cos out, inside is cos^2 +sin^2 , where is 1?

OpenStudy (anonymous):

didn't you say last time that if i take a number out i have to replace it by 1 to show that there was something there?

OpenStudy (anonymous):

the cos that is multiplied by sin^2x

OpenStudy (anonymous):

ok I got what you mean, but 1 sin^2 = sin^2 only.

OpenStudy (anonymous):

I am leaving now, I want to sleep, we can work efficiently tomorrow. I 'll be there at 1:00 for you . Prof Gao asked me to meet to discuss about tutoring . He may ask me be a tutor in lab, but how? with my English? how? no way, right?

OpenStudy (anonymous):

don't worry you will do a great job

OpenStudy (anonymous):

I am understanding math because of you so whatever he says, accept it. you really really deserve it. go to bed we will work on some tomorrow. you can come around 3. it is ok i will do more work in the meantime so we have more to work on

OpenStudy (anonymous):

ok, but I am worried for Sydney and you, if i were a tutor, I cannot say no to other. I don't want sit there and tutor for 016 017 students when Sydney has to take second time around for linear algebra.

OpenStudy (anonymous):

she really need help with both linear and cal 2.

OpenStudy (anonymous):

Summer 1 I take English. Although I let it aside long time but I am big confident to say that it will be back to me after 1 week I focus on. I will review everything from cal2 and try to explain algebra in simplest way to help Sydney

OpenStudy (anonymous):

my delete and backspace are not working. i don't know why

OpenStudy (anonymous):

hey, bthere there is means?

OpenStudy (anonymous):

just close the question is ok, no one come to closed question

OpenStudy (anonymous):

lol by keyboard is attact ed by aliens

OpenStudy (anonymous):

you know that when you say something in fast, i don't understand, now you type in abbreviate, you want me crazy?

OpenStudy (anonymous):

I will see you around 3? take a rest because you and I have a lot of work to do. you will be testing me. A lot

OpenStudy (anonymous):

close the question now

OpenStudy (anonymous):

you should prepare the problems you have, I pick and what if it is too easy or not the problem you don't get. actively working please, only one you know what you need

OpenStudy (anonymous):

I am laughing because i cannot delete anything,.I typed. the keyboard its not working. I felleel like i need anything sometimes other times i feel like i am ready to take the final. Am I making sense?

OpenStudy (anonymous):

I want to sleep, bye you now

OpenStudy (anonymous):

I am going too so i can start early. have a good night see you tomorrow

OpenStudy (anonymous):

thanks ofr evverything